OpenStudy (anonymous):

Evaluate the integral sec(x/7)dx I'm always bad at trigonometric functions T_T

zepdrix (zepdrix):

\[\Large \int\limits \sec\left(\frac{x}{7}\right)\;dx\] Mmm this one is pretty tricky actually :U

OpenStudy (anonymous):

It's hard if you don't quote know the secant integral yet.

zepdrix (zepdrix):

We have to do this little sneaky multiplication (which is really not obvious). You probably wouldn't know to do this unless you've done the problem before. :) So don't worry. We want to multiply the top and bottom by \(\Large \dfrac{\sec\left(\frac{x}{7}\right)+\tan\left(\frac{x}{7}\right)}{\sec\left(\frac{x}{7}\right)+\tan\left(\frac{x}{7}\right)}\)

OpenStudy (psymon):

Yeah, this is an integration that usually gets memorized straightforward.

OpenStudy (anonymous):

Isn't there a way to do it by parts?

zepdrix (zepdrix):

Gives us something like this yes? :U \[\Large \int\limits \frac{\sec^2\left(\frac{x}{7}\right)+\sec\left(\frac{x}{7}\right)\tan\left(\frac{x}{7}\right)}{\sec\left(\frac{x}{7}\right)+\tan\left(\frac{x}{7}\right)}\;dx\] From here it's a simple U-sub.

zepdrix (zepdrix):

By parts? :u hmm i've never tried that before :3

zepdrix (zepdrix):

Following along beau? +_+

OpenStudy (anonymous):

yeah, I'm wondering if this is an only way?

zepdrix (zepdrix):

It's the simplest way, of that I'm pretty certain.

zepdrix (zepdrix):

As wio mentioned before, this is really one that you just want to memorize. If you're being asked to go through the steps though, this is what we need to do. :)

OpenStudy (psymon):

Yeah, without memorizing the integration of it, this is the way you would go about it. For the most part, just try to memorize: \[\int\limits_{}^{}\sec(u)=\ln|\sec(u) +\tan(u)| +C\] Its just in this case your u is x/7. So if you solve for dx, you get du = (1/7)dx and dx = 7du. So this is just a basic secant integral that gets multiplied by 7.

OpenStudy (psymon):

The way I remember the formula is I know the derivative of secant is sec(x)tan(x). So just take that, throw it in an ln and put a plus sign in between.

OpenStudy (anonymous):

ok thank you guys <3

OpenStudy (psymon):

Yep. And before you go, though, this is the same thing youd have to do if it were integral of csc instead. So just to remember that integration: \[\int\limits_{}^{}\csc(u) = -\ln|\cot(u) + \csc(u)| + C\] As for memorizing this one, its the same process. The derivative of csc is -csc(x)cot(x). So just keep the negative outside, throw the csc(x)cot(x) inside of the ln and throw a plus sign in between.

OpenStudy (anonymous):

lol you're my saver <3 csc is the next homework i have to solve xD

OpenStudy (psymon):

Ah xD Lol, yikes. Have you learned how to integrate tangent and cotangent by chance?

OpenStudy (anonymous):

yes i think i'm ok with it now ^_^

OpenStudy (psymon):

Alright, awesome then xD Thats all 6 of the basic trig functions then, so until you come across something funky youre all good then I hope. lol.